OpenStudy (anonymous):

Line BC= 2.7cm, Line Ac= 1.1cm, Angle C = 90 Solve the tirangle

OpenStudy (anonymous):

what you want? the length of BA? or ?

OpenStudy (anonymous):

i need angle a and angle b and line ab

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

since C angle is 90 so it is right angle... BC^2=AC^2+BC^2

OpenStudy (anonymous):

BC^2=(1.1)^2+(2.7)^2 add these numbers and then take the sqr this is BC length

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

about the angles

OpenStudy (anonymous):

using cos or sin or tan function

OpenStudy (anonymous):



OpenStudy (anonymous):

tan(a)=2.7/1.1 a=atan(2.7/1.1) b=180-90-a

OpenStudy (anonymous):

need more help

OpenStudy (anonymous):

no i have it now thank you